OpenStudy (anonymous):

A dam needs a supporting beam. The dam leans at an 80 degree angle and is 200 feet tall. If the base of the supporting beam is place 75 feet from the base of the dam and the beam extends to the top of the dam, how long must the beam be?

OpenStudy (shiraz14):

Using Cosine Rule, h^2 = 200^2 + 75^2 - (2*75*200cos80) h = 201ft (rounded to the nearest integer)

OpenStudy (anonymous):

I am so sorry I had to go to class, I wasn't online when you helped. My bad.

OpenStudy (anonymous):

But thank you!

OpenStudy (shiraz14):

No worries & you're welcome